# Quantum Wave Simulation with Sources and Loss Functions

### v1

@@ 2.2 Maxwell’s Equations
2.3 The Elastic Wave Equation
In the interest of clear notation, we consider the linear elastic wave equation in 3D. Here, the state vector is given by $\hat{\mathbf{w}}_{\text{elastic}}=[\mathbf{v};\text{vec}(\boldsymbol{\sigma})]$, where $\mathbf{v}$ is the particle velocity field and $\text{vec}[\boldsymbol{\sigma}]=[\sigma_{xx};\sigma_{yy};\sigma_{zz};\sigma_{xy};\sigma_{xz};\sigma_{yz}]$ denotes the six independent entries of the positive definite and symmetric stress tensor $\boldsymbol{\sigma}$. The corresponding strain tensor can be represented as $\text{vec}(\boldsymbol{\epsilon})=[\epsilon_{xx};\epsilon_{yy};\epsilon_{zz};2\epsilon_{xy};2\epsilon_{xz};2\epsilon_{yz}]$. This notation allows us to represent the constitutive relations in the Voigt notation [45] as
$\text{vec}[\boldsymbol{\epsilon}](\mathbf{x},t)=\mathbf{S}(\mathbf{x})\text{vec}[\boldsymbol{\sigma}](\mathbf{x},t),$ (16)
where $\mathbf{S}\in\mathbb{R}^{6\times 6}$ is the compliance matrix; a positive definite and invertible matrix for regular elasticity (e.g., odd elasticity can not be covered with this formulation [46]). Note that $\mathbf{S}$ has at most $21$ independent parameters [47]. We can now identify
$\hat{\mathbf{B}}_{\text{elastic}}=\begin{bmatrix}\rho(\mathbf{x})\mathds{1}_{3\times 3}&\mathbf{0}_{3\times 6}\\ \mathbf{0}_{6\times 3}&\mathbf{S}\end{bmatrix}.$ (17)
@@ 2.3 The Elastic Wave Equation
$\hat{\mathbf{D}}=\begin{bmatrix}\partial_{x}&0&0&\partial_{y}&\partial_{z}&0\\ 0&\partial_{y}&0&\partial_{x}&0&\partial_{z}\\ 0&0&\partial_{z}&0&\partial_{x}&\partial_{y}\end{bmatrix}.$ (18)
@@
$\hat{\mathbf{A}}_{\text{elastic}}=\begin{bmatrix}\mathbf{0}_{3\times 6}&\hat{\mathbf{D}}\\ \hat{\mathbf{D}}^{T}&\mathbf{0}_{6\times 3}\end{bmatrix}.$ (19)
Note that $\hat{\mathbf{A}}$ appears symmetric, but it is not: integration by parts, which shifts the derivative in the inner product introduces a minus sign, implying that $(\hat{\mathbf{A}}\hat{\mathbf{w}}|\hat{\mathbf{w}}^{\prime})=-(\hat{\mathbf{w}}|\hat{\mathbf{A}}\hat{\mathbf{w}}^{\prime})$. The elastic energy is now computed as
@@
2.4 Numerical Discretization and Boundary Conditions
@@
Our framework is compatible with any discretization method, provided that the (anti-)Hermitian properties of the operators are preserved. We represent the numerical discretization of the wave field and operators to vectors and matrices as $\hat{\mathbf{w}}\rightarrow\mathbf{w}$, $\hat{\mathbf{H}}\rightarrow\mathbf{H}$, $\hat{\mathbf{B}}\rightarrow\mathbf{B}$, and $\hat{\mathbf{A}}\rightarrow\mathbf{A}$. The discretized equations of motion read
$\frac{d\mathbf{w}_{Q}(t)}{dt}=-i\mathbf{H}\mathbf{w}_{Q}(t)=\mathbf{B}^{-1/2}\mathbf{A}\mathbf{B}^{-1/2}\mathbf{w}_{Q}(t),\quad\mathbf{w}_{Q}(0)=\mathbf{B}^{1/2}\mathbf{w}_{0}.$ (21)
where $\mathbf{A}$ and $\mathbf{B}$ incorporate the BCs (Section C.3).
@@
Preserving the symmetry properties of these operators when discretizing requires that $\mathbf{B}$ is a Hermitian matrix and $\mathbf{A}$ an anti-Hermitian matrix. In Appendix C we exemplify the symmetry-preserving discretization of the acoustic wave equation using finite differences with a staggered grid. We further show that this discretization naturally gives Neumann or Dirichlet BCs on a rectangular domain. Then, starting from the discretization from these particular BCs that respect the symmetries, we discuss in Appendix D how one can introduce a general class of BCs that can be mapped to linear constraints [48]. This includes homogeneous or inhomogeneous Dirichlet or Neumann BCs along non-rectangular domain boundaries.
@@
Additionally, $\mathbf{B}$ must be a diagonal or block-diagonal matrix, with block sizes that are independent of the overall system size; otherwise, $\mathbf{B}^{1/2}$ becomes dense, and the initial quantum state $\mathbf{w}_{Q}(0)$ will lack sparsity. For instance, certain finite element methods produce non-block-diagonal mass matrices. To avoid this issue, we recommend using spectral element methods with Gauss-Lobatto-Legendre integration, which results in diagonal mass matrices [1] or finite differences.
3 Estimating Subspace $l_{2}$-norms and Energies
@@
Reading the complete discrete wave field $\mathbf{w}_{Q}(t)$ comes with exponential and, hence, prohibitive costs. Consequently, it is essential to perform computations with the simulated wave field directly on the QC, e.g., by evaluating loss functions. We show that it is possible to efficiently measure $l_{2}$-norms of subspaces containing multiple superposed quantum states, and that this measurement achieves optimal precision scaling with the number of experiment evaluations:
@@
Let $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\}\in\mathbb{C}^{L}$. Given a projector $\mathbf{P}_{\mathcal{S}}=\text{diag}(p_{1},p_{2},\dots,p_{L})$, which projects onto a subspace $\mathcal{S}\subseteq\mathbb{C}^{L}$, where each $p_{i}\in\{0,1\}$ and $d=\sum_{i=1}^{L}p_{i}$, and let there be a state preparation oracle $\mathcal{U}$ that prepares the quantum state |$\ket{\phi}=[\mathbf{x}_{1};\mathbf{x}_{2};\dots;\mathbf{x}_{M}]/||[\mathbf{x}_{1};\mathbf{x}_{2};\dots;\mathbf{x}_{M}]||$ |; then, if either $d=\mathcal{O}(\text{polylog}(L))$ or $L-d=\mathcal{O}(\text{polylog}(L))$, there exists a quantum algorithm that efficiently estimates the squared sum of $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\}$ on $\mathcal{S}$ as
$l_{\Sigma,\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{x}_{1}+\mathbf{P}_{\mathcal{S}}\mathbf{x}_{2}+\dots+\mathbf{P}_{\mathcal{S}}\mathbf{x}_{M}||^{2},$ (22)
@@
We will now apply this result to compare two wave fields, such as a simulated and a target $\mathbf{B}^{1/2}$-transformed wave field and to measure subvolume energies (see Section 2). By setting $M=2,\;\mathbf{x}_{1}=\mathbf{w}_{Q},\;\mathbf{x}_{2}=-\mathbf{w}_{Q}^{\text{target}}$, where $\mathbf{w}_{Q}$ is the simulated wave field and $\mathbf{w}_{Q}^{\text{target}}$ is a target or observed wave field which is only defined or known within subvolume $\mathcal{S}$, we can write
$l_{\Sigma,\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{w}_{Q}-\mathbf{P}_{\mathcal{S}}\mathbf{w}_{Q}^{\text{target}}||^{2}.$ (23)
@@
This corresponds to an estimate of the $l_{2}$-distance between both fields on subspace $\mathcal{S}$. By setting $M=1,\;\mathbf{x}_{1}=\mathbf{w}_{Q}$ we obtain
$l_{\Sigma,\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{w}_{Q}||^{2},$ (24)
@@
which estimates the energy of $\mathbf{w}$ on $\mathcal{S}$. Theorem 2 can also be used to compute energies and $l_{2}$-comparisons of a superposition of multiple wave fields, extending the work of [35] on measuring energies. These multi-state $l_{2}$- and energy estimates are key components of the source implementation algorithm, which we will introduce in Section 4.
Corollary 3 (Weighted $l_{2}$-norm).
$l_{\Sigma,v}^{2}=\sum_{j}^{V}v_{j}l_{\Sigma,\mathcal{S}_{j}}^{2},$ (25)
Applying this to the comparison between two wave fields we obtain
$l_{\Sigma,B}^{2}=\sum_{j}^{\bar{B}}B^{-1/2}_{j}||\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}_{Q}-\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}_{Q}^{\text{target}}||^{2}=\sum_{j}^{\bar{B}}||\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}-\mathbf{P}_{\mathcal{S}_{j}}\mathbf{w}^{\text{target}}||^{2},$ (26)
where $\bar{B}$ is the number of unique weights in the diagonal matrix $\mathbf{B}$, i.e. unique material parameters, and $S_{j}$ is the subspace only containing grid points with material parameters $B_{j}$. Here, we need to assume that $\mathbf{B}$ is diagonal in $\mathcal{S}$, which holds true for the acoustic wave equation or the isotropic Maxwell’s equation.
@@
Given an efficient quantum algorithm for a unitary basis transformation, comparisons can also be performed in bases other than the computational basis. This, for example, includes the spatial frequency domain, which can be efficiently accessed through the quantum Fourier transform.
4 Implementing Sources
@@
Incorporating time-dependent sources into the HS of classical wave equations poses a significant challenge when aiming to maintain efficient runtime scaling. We consider source terms of the form
$\hat{\mathbf{s}}(\mathbf{x},t)=\sum_{s=1}^{S}\boldsymbol{\chi}_{s}(\mathbf{x})f_{s}(t),$ (27)
where $\boldsymbol{\chi}_{s}(\mathbf{x})\in\mathbb{H}^{L}$ represents the spatial distribution and $f_{s}(t)$ the temporal function of the $s^{\text{th}}$ source. We address the case of $s$ point sources in space, i.e., $\boldsymbol{\chi}_{s}(\mathbf{x})=\hat{\boldsymbol{\chi}}\delta(\mathbf{x}-\mathbf{x}_{s})$, where $\hat{\boldsymbol{\chi}}\in\mathbb{C}^{L}$ is any vector.
@@
Assuming that $f_{s}(t)$ is a pulse function, whose support in time is compact and with a length either constant or inversely proportional to its the maximum frequency, we will show that we can classically map the source to sparse initial conditions with an implementation complexity of $\mathcal{O}(1)$. This already covers many practically relevant source scenarios. We then extend our approach to a broader class of sources that lack compact support in time.
4.1 Single Pulse Source

### v3

@@ II.2 Maxwell’s Equations
[…]
$=2E_{\text{electric}}(t)+2E_{\text{magnetic}}(t).$
II.3 The Elastic Wave Equation
In the interest a of clear notation, we consider the linear elastic wave equation in 3D. Here, the state vector is given by $\hat{\mathbf{w}}_{\text{elastic}}=[\hat{\mathbf{v}};\text{vec}(\boldsymbol{\hat{\sigma}})]\in\{f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{9}\}$, where $\hat{\mathbf{v}}$ is the particle velocity field and $\text{vec}[\boldsymbol{\hat{\sigma}}]=[\sigma_{xx};\sigma_{yy};\sigma_{zz};\sigma_{xy};\sigma_{xz};\sigma_{yz}]$ denotes the six independent entries of the positive definite and symmetric stress tensor $\hat{\boldsymbol{\sigma}}$. The corresponding strain tensor can be represented as $\text{vec}(\boldsymbol{\hat{\epsilon}})=[\epsilon_{xx};\epsilon_{yy};\epsilon_{zz};2\epsilon_{xy};2\epsilon_{xz};2\epsilon_{yz}]$. This notation allows us to represent the constitutive relations in the Voigt notation as [48]
$\text{vec}[\boldsymbol{\hat{\epsilon}}](\mathbf{x},t)=\mathbf{S}(\mathbf{x})\text{vec}[\boldsymbol{\hat{\sigma}}](\mathbf{x},t),$ (16)
where $\mathbf{S}\in\{f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{6\times 6}\}$ is the compliance matrix; a positive definite and invertible matrix for regular elasticity (e.g., odd elasticity [49] cannot be covered with this formulation). Note that $\mathbf{S}$ has at most $21$ independent parameters [50]. We can now identify
$\hat{\mathbf{B}}_{\text{elastic}}(\mathbf{x})=\begin{bmatrix}\rho(\mathbf{x})\mathds{1}_{3\times 3}&\mathbf{0}_{3\times 6}\\ \mathbf{0}_{6\times 3}&\mathbf{S}\end{bmatrix}.$ (17)
@@ II.3 The Elastic Wave Equation
$\mathcal{D}=\begin{bmatrix}\partial_{x}&0&0&\partial_{y}&\partial_{z}&0\\ 0&\partial_{y}&0&\partial_{x}&0&\partial_{z}\\ 0&0&\partial_{z}&0&\partial_{x}&\partial_{y}\end{bmatrix}.$ (18)
@@
$\hat{\mathbf{A}}_{\text{elastic}}=\begin{bmatrix}\mathbf{0}_{3\times 6}&\mathcal{D}\\ \mathcal{D}^{T}&\mathbf{0}_{6\times 3}\end{bmatrix}.$ (19)
Note that $\hat{\mathbf{A}}$ appears symmetric, but it is not: integration by parts, which shifts the derivative in the inner product introduces a minus sign, implying that $(\hat{\mathbf{A}}\hat{\mathbf{w}}|\hat{\mathbf{w}}^{\prime})=-(\hat{\mathbf{w}}|\hat{\mathbf{A}}\hat{\mathbf{w}}^{\prime})$. The elastic energy is computed as [51]
@@
$=2E_{\text{kinetic}}(t)+2E_{\text{potential}}(t).$
II.4 Numerical Discretization and Boundary Conditions
@@
Our framework is compatible with any discretization method, provided that the (anti-)Hermitian properties of the operators are preserved. We represent the numerical discretization of the wave field and operators to vectors and matrices as $\hat{\mathbf{w}}\rightarrow\mathbf{w}$, $\hat{\mathbf{H}}\rightarrow\mathbf{H}$, $\hat{\mathbf{B}}\rightarrow\mathbf{B}$, and $\hat{\mathbf{A}}\rightarrow\mathbf{A}$, where $\mathbf{w}\in\mathbb{R}^{L}$. The discretized equations of motion read
$\frac{d\mathbf{w}_{Q}(t)}{dt}$ $=-i\mathbf{H}\mathbf{w}_{Q}(t)=\mathbf{B}^{-1/2}\mathbf{A}\mathbf{B}^{-1/2}\mathbf{w}_{Q}(t),$ (21)
$\mathbf{w}_{Q}(0)$ $=\mathbf{B}^{1/2}\mathbf{w}_{0}.$
where $\mathbf{A}$ and $\mathbf{B}$ incorporate the BC (Section C.3).
@@
Preserving the symmetry properties of these operators when discretizing requires that $\mathbf{B}$ is a Hermitian matrix and $\mathbf{A}$ an anti-Hermitian matrix. In Appendix C we exemplify the symmetry-preserving discretization of the acoustic wave equation using finite differences with a staggered grid. We further show that this discretization naturally gives Neumann or Dirichlet BC on a rectangular domain. Then, starting from the discretization of these particular BC that respect the symmetries, we discuss in Appendix D how one can introduce a general class of BC that can be mapped to linear constraints [52]. This includes homogeneous or inhomogeneous Dirichlet or Neumann BC along non-rectangular domain boundaries.
@@
Additionally, $\mathbf{B}$ must be a diagonal or block-diagonal matrix, with block sizes that are independent of the overall system size; otherwise, $\mathbf{B}^{1/2}$ becomes dense, and the initial quantum state $\mathbf{w}_{Q}(0)$ will lack sparsity.
III Runtime Advantage
Given that the discretized Hamiltonian $\mathbf{H}$ (8) has at most $d=\mathcal{O}(1)$ nonzero values in each row and column, which is ensured when $\mathbf{A}$ and $\mathbf{B}$ are $d$-sparse matrices, and if neither the bit precision $m$ nor the error $\epsilon$ is dominating, quantum simulation requires $\mathcal{O}(nt)=\mathcal{O}(\log Nt)$ operations (Theorem 1). In principle this offers an exponential speed-up over classical solvers, which all scale at least with $\mathcal{O}(Nt)$ [53, 54, 55, 56, 57]. However, as we now show, the obtainable speed-up is usually only polynomial [36]. In wave propagation problems, increasing grid points typically aims to (i) enlarge the physical domain or (ii) increase the simulated frequencies. For systems with local couplings, such as those resulting from discretizing the considered wave equations, enlarging the domain requires a longer simulation time $t$ for waves to reach the boundaries. Similarly, a finer discretization to simulate higher frequencies raises $\|\mathbf{H}\|_{\max}$ due to the reduced grid spacing. As per Theorem 1, the number of query calls to $\mathcal{H}$ scales linearly with both $t$ and $\|\mathbf{H}\|_{\max}$. Consequently, the quantum speed-up over classical algorithms becomes a polynomial factor dependent on the spatial dimension $D$. Specifically, let $t=\mathcal{O}(N^{1/D})$ represent the simulation time for the signal to reach the domain boundary in $D$ dimensions. For classical algorithms scaling linearly with $N$, the overall runtime complexity is $\mathcal{O}(Nt)=\mathcal{O}(N^{1+1/D})$. In contrast, the HS algorithm has runtime $\tilde{\mathcal{O}}(t)=\tilde{\mathcal{O}}(N^{1/D})$, where the tilde hides logarithmic factors. Thus, we expect a quadratic speed-up in 1D, cubic in 2D, and quartic in 3D. In contrast, [36] demonstrate an exponential speed-up for oscillator models with purely non-local couplings, where the simulation time does not scale with the system size.
Additionally, the algorithm’s oracle $\mathcal{H}$ imposes constraints: to preserve quantum speed-up, the information transferred to the QC must be of order $\tilde{\mathcal{O}}(N^{1/D})$. This limits the number of independent material parameters to this order, necessitating functional or low-dimensional material representations.
IV Estimating Subspace $l_{2}$-norms and Energies
@@
Reading the complete discrete wave field $\mathbf{w}_{Q}(t)$ from a quantum state comes with exponential and, hence, prohibitive costs. Consequently, it is essential to perform computations with the simulated wave field directly on the QC, e.g., by evaluating loss functions. We show that it is possible to efficiently measure $l_{2}$-norms of state subspaces containing multiple superposed quantum states, and that this measurement achieves optimal precision scaling with the number of experiment evaluations:
@@
Let $\{\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{M}\}\in\mathbb{C}^{L}$. Given a projector $\mathbf{P}_{\mathcal{S}}=\text{diag}(p_{1},p_{2},\dots,p_{L})$, which projects onto a subspace $\mathcal{S}\subseteq\mathbb{C}^{L}$, where each $p_{i}\in\{0,1\}$ and $d=\sum_{i=1}^{L}p_{i}$. Let there be a state preparation oracle $\mathcal{U}$ that prepares the quantum state |$\ket{\phi}=[\mathbf{w}_{1};\mathbf{w}_{2};\dots;\mathbf{w}_{M}]/||[\mathbf{w}_{1};\mathbf{w}_{2};\dots;\mathbf{w}_{M}]||$ |. Then, if either $d=\mathcal{O}(\text{polylog}(L))$ or $L-d=\mathcal{O}(\text{polylog}(L))$, there exists a quantum algorithm that efficiently estimates the squared sum of $\{\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{M}\}$ on $\mathcal{S}$ as
$l_{\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{w}_{1}+\mathbf{P}_{\mathcal{S}}\mathbf{w}_{2}+\dots+\mathbf{P}_{\mathcal{S}}\mathbf{w}_{M}||^{2},$ (22)
@@
We will now apply this result to compare two wave fields, such as a simulated and a target $\mathbf{B}^{1/2}$-transformed wave field and to measure subvolume energies (see Section II). By setting $M=2,\;\mathbf{w}_{1}=\mathbf{w}_{Q},\;\mathbf{w}_{2}=-\mathbf{w}_{Q}^{\text{target}}$, where $\mathbf{w}_{Q}$ is the simulated wave field and $\mathbf{w}_{Q}^{\text{target}}$ is a target or observed wave field which is only defined or known within subvolume $\mathcal{S}$, we can write
$l_{\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{w}_{Q}-\mathbf{P}_{\mathcal{S}}\mathbf{w}_{Q}^{\text{target}}||^{2}.$ (23)
@@
This corresponds to an estimate of the $l_{2}$-distance between both fields on subspace $\mathcal{S}$. By setting $M=1,\;\mathbf{w}_{1}=\mathbf{w}_{Q}$ we obtain
$l_{\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{w}_{Q}||^{2},$ (24)
@@
which estimates the energy of $\mathbf{w}$ on $\mathcal{S}$. Theorem 2 can also be used to compute energies and $l_{2}$-comparisons of a superposition of multiple wave fields, extending the work of [36] on measuring energies. These multi-state $l_{2}$- and energy estimates are key components of the source implementation algorithm, which we will introduce in Section V.
We can utilize a weighted inner product to compare states in the standard, non-$\mathbf{B}^{1/2}$-transformed bases, as detailed in Section A.3. However, as we discuss in Section VI.1 this will result in globality issues in general. In practice, e.g., within the domain of inverse problems, it is usually sufficient to compare transformed wave fields.
@@
Finally, given an efficient quantum algorithm for a unitary basis transformation, comparisons can also be performed in bases other than the computational basis. This, for example, includes the spatial frequency domain, which can be efficiently accessed through the quantum Fourier transform [22].
V Implementing Sources
@@
Incorporating time-dependent sources into the HS of classical wave equations poses a significant challenge when aiming to maintain efficient runtime scaling, as in this framework the wave evolution is unitary and, hence, inherently source free. We consider a source term that is constituted of $S$ sources,
$\hat{\mathbf{s}}(\mathbf{x},t)=\sum_{s=1}^{S}\pmb{\chi}_{s}(\mathbf{x})f_{s}(t),$ (25)
where $\pmb{\chi}_{s}\in\mathbb{R}^{D}\rightarrow\mathbb{R}^{C}$ is the spatial distribution of the $s$-th sources and $f_{s}(t)$ is its source time funcion.
@@
Assuming that $\pmb{\chi}_{s}$ is a point source and $f_{s}$ is a pulse function, whose support in time is compact and with a length either constant or inversely proportional to the maximum frequency, we will show that we can classically map the source to sparse initial conditions with an implementation complexity of $\mathcal{O}(1)$. This already covers many practically relevant source scenarios. We then show how to implement sources that lack compact support in time, but that give rise to rotationally covariant wave fields.
V.1 Single Pulse Source
Consider a single point-source defined by
$\hat{\mathbf{s}}(\mathbf{x},t)=\hat{\pmb{\chi}}\delta(\mathbf{x}-\mathbf{x}_{s})f(t),$ (26)